OpenStudy (anonymous):

So here's the deal. I'll put the logarithm bases into (). Solve log(cosx)sinx+ log(sinx)cosx=2.

OpenStudy (anonymous):

Whew, the bases are functions of x? That's pretty gnarly. Mixing logs and trig functions . . . will probably end up with complex numbers. This will take some work.

OpenStudy (anonymous):

It should be done without complex numbers.

OpenStudy (anonymous):

My best guess is pi/4.

OpenStudy (anonymous):

There may be other solutions, but I know that works.

OpenStudy (anonymous):

Want to know how I did it?

OpenStudy (anonymous):

Sure.

OpenStudy (anonymous):

I took the simplest case in which \[\log_{\cos x}(\sin x) =\log_{\sin x}(\cos x) = 1.\] A logarithm equals 1 when the base and what it is operating on are equal. sin x = cos x at pi/4.

OpenStudy (anonymous):

That's true.

OpenStudy (anonymous):

It'll take a little more doing to see if there are other solutions.

OpenStudy (anonymous):

e.g. log_{cos x}(sin x) could be 1.5 and log_{sin x}(cos x) could be 0.5 and so on.

OpenStudy (anonymous):

That's also true.

OpenStudy (anonymous):

I think I can continue from now on. You gave me the ideas I needed. Thank you very much!

OpenStudy (anonymous):

Thank you. That was a fun one!